We dissolve 2.77 g of an unknown acid, HA, in enough water to produce 25.0 mL of solution. The pH of this solution of HA(aq) is 1.33. We titrate this solution with a 0.250 M solution of NaOH. It takes 41.4 mL of the NaOH solution to reach the equivalence point.

(a) What is the molar mass of HA?

(b) What is the pKa value of HA(aq)?

(c) What is the pH at the equivalence point?

(d) What is the pH one third along the titration when you still have twice as much HA(aq) as A-(aq)?
Answer #1

a) MY MLV2 Mi = molarity of HA M2 = molarity of Maon 1225ml V₂ = 41.4ml M, 8 2520 25xury M, 20.414M Alo. Molarity of onols voPy = -log [H3O+ 1.33 = -log [1430+] (H30] = 10^4:33 [420t] =0.04677 ..Kg = 5.95810-3 pka = -log ka 2.23 e) RM at equivalence